How much clay do you need to add to 4 cubic yards of soil with 40% clay in order to make a soil with 60% clay?
Answer by Boreal(15235)

x=amount of clay
(4)(0.40)+x=(4+x)(0.60), the amount of pure clay in 4 yds^3 at 40% is 1.6 yds^3. We add x amount of pure 100% clay
1.6+x=2.4+0.60x
-0.8=-0.4x
x=2 yds^3 ANSWER
check
2/3 is 40%, 1/3 is 100%
40+40+100 all divided by 3 is 60